OpenStudy (tester97):

How do I use the quadratic formula to solve this equation 4x^2-6x-10=0

OpenStudy (anonymous):

4x^2-6x-10=0 is like: Ax^2+Bx+C=0 A=4 B=-6 C=-10 \[\frac{ -B \pm \sqrt{B^2-4AC} }{ 2A } = X\]

OpenStudy (tester97):

so i just plug in the rite numbers in da quad form?

OpenStudy (anonymous):

yep ^_^ you will do the equation twice, because there is the plus minus symbol. so the first time you can do plus and the second time you can do minus

OpenStudy (tester97):

ok let me solve this and then can u check it?

OpenStudy (anonymous):

yep ^_^

OpenStudy (tester97):

is it -1 and/or 5/2?

OpenStudy (anonymous):

good job ^_^ -1 and 5/2 is correct

OpenStudy (tester97):

thx :D